OpenStudy (anonymous):

Find an explicit formula for the sequence defined recursively by \[a _{0}=7\] \[a _{1}=10\] \[a _{k}=2a _{k-1}-a _{k-2}\]

OpenStudy (mr.math):

\[a(n)=3n+7\].

OpenStudy (anonymous):

Your sort of guess for these right

OpenStudy (mr.math):

\[a_0=7\] \[a_1=3+7\] \[a_2=2(10)-7=20-7=2(3)+7.\]

OpenStudy (mr.math):

Find several \(a_n\)'s and try to see the pattern, Define a formula based on that pattern, and then prove it's true for all n in your domain.

OpenStudy (anonymous):

Can we prove with induction

OpenStudy (mr.math):

I think so.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

so up to a_4

OpenStudy (anonymous):

my sequence from a_0 is 7,10,20,30,40

OpenStudy (mr.math):

How?

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

totally forgot to subtract 7 in a_2

OpenStudy (mr.math):

\[a_0=7\] \[a_1=10\] \[a_2=20-7=13\] \[a_3=26-10=16\] \[a_4=32-13=19\] ....

OpenStudy (anonymous):

ah ok

OpenStudy (mr.math):

You can see the pattern, right?

OpenStudy (anonymous):

not sure if i would have been able to spot it as fast as you did, but i see it now yea

OpenStudy (mr.math):

I think you can, maybe a couple of minutes slower :P At least you can spot it's a linear relation, and this is the easiest type.

OpenStudy (anonymous):

yea, nonlinear are harder

OpenStudy (across):

\[f(n)=3n+7,\]\[n\in\mathbb{N},\]\[\mathbb{N}=\left\{0,1,2,...\right\}.\] Proof: For n = 0,\[f(0)=3(0)+7=7\]which is true. Then assume it is true for n = k and prove that it is true for k+1.\[f(k+1)=f(k)+3,\]\[f(k+1)=3(k+1)+7=3k+3+7=3k+10=(3k+7)+3=f(k)+3.\blacklozenge\]yup